OpenStudy (anonymous):

what is the exact value of the square root of 28 over 9 square root of 16

OpenStudy (2bornot2b):

Do you mean \[\frac{\sqrt{28}}{9}\sqrt{16}\] Or do you mean \[\frac{\sqrt{28}}{9\sqrt{16}}\]

OpenStudy (anonymous):

sqrt 28 = sqrt4 * sqrt 7 = 2sqrt7 sqrt 16 = 4 so your expression 2sqrt7 = sqrt7 / 18 ----- 9* 4